5. An object traveling with a speed of 10 m/s slides on a horizontal floor. How far will it travel before coming to rest if the coefficient of friction is 0.30?


1
Expert's answer
2021-09-21T13:57:24-0400

"F_f=ma\to \mu N=ma\to\mu mg=ma\to a=\mu g=0.3\cdot9.81=2.943\ (m/s^2)"


"s=\frac{v^2}{2a}=\frac{10^2}{2\cdot2.943}\approx17 \ (m)" . Answer